make a linear equation base on the problem:
greg, who is 2 years older than liam, is celebrating his 9th birthday. how old would both boys be after 3 years? after 7 years? after 10 years?

Answer by Theo(13342)

greg is 2 years older than liam.
greg is celebrating his 9th birthday.
how old would both boys be after 3, 7, 10 years.

when greg is 9, liam is 7
in 3 years, greg will be 12 and liam will be 10.
in 7 years, greg will be 16 and liam will be 14.
in 10 years, greg will be 19 and liam will be 17.

to make an equation out of this, let x = liam's age and let y = greg's age.

since greg is 2 years older than liam, you get:

y = x + 2

x represents liam's age and y represents greg's age.

when x = 7, y = 9
when x = 7 + 3 = 10, y = 12
when x = 7 + 7 = 14, y = 16
when x = 7 + 10 = 17, y = 19

the equation can be graphed as shown below: